OpenStudy (anonymous):

How do I do this and what is the answer? Given the equation y - 4 = 3/4 (x + 8) in point-slope form, identify the equation of the same line in standard form. A) -3/4 x + y = 10 B) 3x - 4y = -40 C) y = 3/4 x + 12 D) y = 3/4 x + 10

OpenStudy (anonymous):

point slope form is : y-y1= m (x-x1) so y=4 m= 3/4 x=8 standard form is ax+by=c so distribute m in the original equation y - 4 = (3/4x + 6) move -4 to the other side by adding y= 3/4x +10 re-arrange this information into standard ax+by=c form -3/4x+y=10